Please just don't get it....Suppose that the function?
Suppose that the function P=13+45ln x represents the percentage of inbound e-mail in the U.S. that is consideed spam, where x is the number of years after 2000.
Carry all calculations to six decimals on each intermediate step, when necessary.
a) Use this model to approximate the percentage of spam in the year 2003.
Answer:
b) Use this model to approximate the year that the percent of spam will reach 95% provided that law enforcement regarding spammers does not change?
Answer:

Answer by stanbon(75887)

Suppose that the function P=13+45ln x represents the percentage of inbound e-mail in the U.S. that is consideed spam, where x is the number of years after 2000.
Carry all calculations to six decimals on each intermediate step, when necessary.
a) Use this model to approximate the percentage of spam in the year 2003.
Answer: P(2003-2000) = 13+45ln(3) = 62.4376%
----------------------
b) Use this model to approximate the year that the percent of spam will reach 95% provided that law enforcement regarding spammers does not change?
95 = 13 + 45ln(x)
82 = 45ln(x)
ln(x) = 1.8222
x = e^1.8222 = 6.1856
------------------
95% spam will be reached in 2000+6.1856 or in the year 2007
==================
Cheers,
Stan H.